# Improving Robustness Estimates in Natural Language Explainable AI though Synonymity Weighted Similarity Measures

### v1

@@ 5.2 Experimental Data
To test the effects of synonymity weighting on stability estimates, we require examples generated from an adversarial XAI process. We use the method in Burger et al. (2023) that easily allows one to replace the similarity measure used to guide the search process. As the similarity measure used for the comparison of explanations controls the search, simply comparing the end result calculated under a single measure will not accurately describe the stability, as the replacement of words with viable perturbations can differ appreciably between measures used. We note that the attack process itself is unmodified in that synonymity weighting is not applied to the measure guiding the attack but only to the final output. This was done to prevent unnecessary computation as repeating the attack with the added synonymity weighting would have required over a month of additional computation on an A6000 GPU. This renders our conclusions understated as our construction of the synonymity weighting results, at worst, in the same similarity to the original measure. For all practical purposes, the similarity will increase, which will reduce the attack success rate.
To create the raw material for our analysis, we generate batches of 50 adversarial examples using the algorithm in Burger et al. (2023). Each batch is generated with respect to a similarity measure and a success threshold. The four measures defined in Section 5.1 are used: The Jaccard index, Kendall’s Tau Rank Distance, and Spearman’s footrule, and RBO with weighting parameters 0.5, 0.7, and 0.9. Combined with the thresholds of 30%, 40%, 50%, and 60%, this results in 1,200 adversarial examples per data set. For the data sets, we use two of those included in Burger et al. (2023) and their associated pre-trained DistilBERT models. The first data set is the Twitter dataset with short length (average 11 words) gender bias Twitter dataset (GB) and the second is the data set of symptoms to diagnosis of moderate length (average 29 words) symptoms to diagnosis dataset (S2D).
@@ 6 Results & Discussion
Jaccard & Spearman: Immediately seen is the drastic reduction in the attack success rate for the Jaccard index and Spearman’s footrule. For Jaccard under the GB dataset, we see the three most stringent thresholds (30%, 40%, and 50%) reduced to zero successes while the 60% threshold is reduced to 5%. Similar results hold for the S2D dataset. Spearman is even more sensitive than Jaccard, with every S2D success under the original calculation changed to failure. Spearman’s footrule was investigated to determine if the weighting scheme was appropriate due to the enormous success reduction and comparatively more complex weighting formulation over Jaccard. However, the different choices of the penalty value showed little change in the results with only slight ($\leq 5\%$) reductions in calculated similarity. These results demonstrate that both standard measures are extremely sensitive to changes in the explanation that are due to properties of the measure, and not due to a change in fundamental meaning of the explanation. With such significant decreases in attack success, any conclusions generated about XAI stability under these measures should be viewed with suspicion.
RBO: In contrast to Jaccard and Spearman, RBO gains much less from synonymity weighting, only showing minor changes in success rate for weighting values of 0.5 and 0.7. For weighting value 0.9 it is not surprising to see more efficacy as a more uniform distribution of importance to each feature. There still exist important examples where RBO can profit through the use of synonymity weighting (Table 1), but in general the changes are modest (Table 2). The lack of overall success associated with RBO0.9 makes a firm judgement on its usefulness premature, data with more successful attacks is required. The appreciable difference in success rate change between RBO and Jaccard was not expected as both are intersection-based. RBO’s relatively minor gain from synonymity weighting is likely due to its own inherent weighting scheme; to confirm this, the intrinsically weighted version of the other measures can be tested, and we leave this for future investigation. Overall, RBO remains a strong candidate for use in adversarial XAI as its design innately maintains a balance between sensitivity and indifference.
@@ 7 Limitations and Conclusion
The constructions for the specific synonymity weighting definitions may not be optimal and are intended largely as a demonstration of the concept. Spearman’s footrule may benefit from a new method that may prevent the severe change in similarity between the original and weighted versions. Our choice of numerical synonymity is another prime choice for optimization, especially as part of speech checking is not associated with our selection. Superior estimates of synonymity can probably be obtained with alternative methods.
Overall, this work demonstrates the usefulness of synonymity weighting to allow superior estimates of XAI stability. Substantial reductions in attack success rate for certain measures show the possibility of extremely understating XAI stability without the judicious selection of an appropriate adversarial perturbation method. The application of this weighting scheme also incurs negligible computational overhead as the bottleneck in current algorithms is the explanation generation itself, so even measures who see only moderate impact from synonymity weighting can incorporate it without burden.

### v2

@@ 5.2 Experimental Data
To test the effects of synonymity weighting on stability estimates, we require examples generated from an adversarial XAI process. We use the method in [burger2023explanations] that easily allows one to replace the similarity measure used to guide the search process. As the similarity measure used for the comparison of explanations controls the search, simply comparing the end result calculated under a single measure will not accurately describe the stability, as the replacement of words with viable perturbations can differ appreciably between measures used.
We note that the attack process itself is unmodified in that synonymity weighting is not applied to the measure guiding the attack but only to the final output. This was done to prevent unnecessary computation as repeating the attack with the added synonymity weighting would have required over a month of additional computation on an A6000 GPU. This renders our conclusions understated, as our construction of the synonymity weighting results, at worst, in the same similarity to the original measure. For all practical purposes, the similarity will increase, which will reduce the attack success rate.
To create the raw material for our analysis, we generate batches of 50 adversarial examples using the algorithm in [burger2023explanations]. Each batch is generated with respect to a similarity measure and a success threshold. The four measures defined in Section 5.1 are used: The Jaccard index, Kendall’s Tau Rank Distance, and Spearman’s footrule, and RBO with weighting parameters 0.5, 0.7, and 0.9. Combined with the thresholds of 30%, 40%, 50%, and 60%, this results in 1,200 adversarial examples per data set. For the data sets, we use two of those included in [burger2023explanations] and their associated pre-trained DistilBERT models. The first data set is the Twitter dataset with short length (average 11 words) gender bias Twitter dataset (GB) and the second is the data set of symptoms to diagnosis of moderate length (average 29 words) symptoms to diagnosis dataset (S2D).
@@
Table 5: Comparative average success rates across methods for measuring similarity
Jaccard Kendall Spearman RBO0.5 RBO0.7 RBO0.9
Base Synw Base Synw Base Synw Base Synw Base Synw Base Synw
GB GloVe (Original) 0.54 0.01 0.98 0.88 0.59 0.03 0.30 0.27 0.19 0.13 0.09 0.00
fastText 0.54 0.02 0.98 0.88 0.59 0.04 0.30 0.27 0.19 0.13 0.09 0.01
WordNet 0.54 0.34 0.98 0.93 0.59 0.00 0.30 0.29 0.19 0.18 0.09 0.05
S2D GloVe (Original) 0.64 0.08 1.00 0.96 0.67 0.02 0.18 0.17 0.14 0.11 0.09 0.03
fastText 0.64 0.07 1.00 0.95 0.67 0.02 0.18 0.17 0.14 0.12 0.09 0.03
WordNet 0.64 0.28 1.00 0.98 0.67 0.01 0.18 0.17 0.14 0.12 0.09 0.08
Table 6: Comparative average similarity levels across methods for measuring similarity
Jaccard Kendall Spearman RBO0.5 RBO0.7 RBO0.9
Base Synw Base Synw Base Synw Base Synw Base Synw Base Synw
GB GloVe (Original) 0.42 0.59 0.28 0.30 0.39 0.61 0.35 0.36 0.34 0.36 0.26 0.31
fastText 0.42 0.57 0.28 0.30 0.39 0.61 0.35 0.36 0.34 0.36 0.26 0.30
WordNet 0.42 0.51 0.28 0.29 0.39 0.79 0.35 0.35 0.34 0.35 0.26 0.28
S2D GloVe (Original) 0.43 0.54 0.31 0.32 0.41 0.62 0.37 0.37 0.43 0.44 0.27 0.30
fastText 0.43 0.53 0.31 0.32 0.41 0.63 0.37 0.37 0.43 0.43 0.27 0.29
WordNet 0.43 0.48 0.31 0.32 0.41 0.71 0.37 0.37 0.43 0.43 0.27 0.28
@@ 6 Results & Discussion
Jaccard & Spearman: Immediately seen is the drastic reduction in the attack success rate for the Jaccard index and Spearman’s footrule. For Jaccard under the GB dataset, we see success rates for the GB dataset under Jaccard fall to nearly zero for all but the most lenient (60%) threshold. Similar results hold for the S2D dataset. Spearman is even more sensitive than Jaccard, with every S2D success under the original calculation changed to failure. Spearman’s footrule was investigated to determine if the weighting scheme was appropriate due to the enormous success reduction and comparatively more complex weighting formulation over Jaccard. However, the different choices of the penalty value showed little change in the results with only slight ($\leq 5\%$) reductions in calculated similarity. These results demonstrate that both standard measures are extremely sensitive to changes in the explanation that are due to properties of the measure, and not due to a change in fundamental meaning of the explanation. With such significant decreases in attack success, any conclusions generated about XAI stability under these measures should be viewed with suspicion.
RBO: In contrast to Jaccard and Spearman, RBO gains less from synonymity weighting, only showing minor changes in success rate for weighting values of 0.5 and 0.7. For weighting value 0.9 it is not surprising to see more efficacy as a more uniform distribution of importance to each feature. There still exist important examples where RBO can profit through the use of synonymity weighting (Table 1), but in general the changes are modest (Table 2). The lack of overall success associated with RBO0.9 makes a judgement on its usefulness premature without more data from successful attacks. The appreciable difference in success rate change between RBO and Jaccard was not expected as both are intersection-based. RBO’s relatively minor gain from synonymity weighting is likely due to its own inherent weighting scheme; to confirm this, the intrinsically weighted version of the other measures can be tested, and we leave this for future investigation. Overall, RBO remains a strong candidate for use in adversarial XAI as its design innately maintains a balance between sensitivity and indifference.
@@
6.1 Sensitivity Analysis
To provide robustness in our conclusions about synonymity weighting we performed a sensitivity analysis that adds two methodologically different options: fastText, which leverages subword information, and a thesaurus-based method derived from WordNet. This analysis was conducted with respect to average success and similarity levels across the two datasets (Tables 5-6).
Noticeably, there is little difference between the original GloVe embedding used and fastText, which is generally considered to be a superior alternative. The only deviations between GloVe and fastText in the success rate analysis is 1%, indicating that both embeddings are largely interchangeable under our method despite their different construction. The thesaurus-based option using WordNet showed more difference as expected due to its dichotomous construction where words are either synonyms and so are given full value (1) or not and given a value of (0). In general, the thesaurus-based method produced values closer to those calculated without synonymity weighting over the embedding based methods. However, most values were fairly close with only the Jaccard measure providing noteworthy mention. The success rates were substantially higher ($\sim 32\%$ for GB and $\sim 21\%$ S2D) compared to the embedding-based methods, but the overall success rate compared to the original base was far lower with a 20% reduction on GB and a 36% reduction on S2D. Overall, the choice of measure for generating the word-to-word similarity values does not appear to be particularly important and all choices tested result in superior estimates of XAI stability.
@@ 7 Limitations and Conclusion
The weighting constructions presented are intended as a proof-of-concept and may not be optimal. A limitation is that our analysis applies synonymity weighting post hoc; future work could integrate these weighted measures directly into the adversarial search process. This could be particularly beneficial for measures like Spearman’s footrule, where a more integrated approach might mitigate the severe changes in similarity observed. Furthermore, while our choice of numerical synonymity is robust, superior estimates could be obtained using more advanced techniques, such as contextual embeddings from transformer models like BERT or by incorporating part-of-speech data. Finally, beyond improving the measures, future work could apply this robust evaluation framework to benchmark the stability of different XAI methods against each other, providing a fairer comparison of their stability.
Overall, this work demonstrates the usefulness of synonymity weighting to allow superior estimates of XAI stability. Substantial reductions in attack success rate for certain measures show the possibility of extremely understating XAI stability without the judicious selection of an appropriate adversarial perturbation method. The application of this weighting scheme also incurs negligible computational overhead as the bottleneck in current algorithms is the explanation generation itself, so even measures that see only moderate impact from synonymity weighting can incorporate it without burden. With the inclusion of synonymity in these measures, we offer a more robust method for threat detection in AI systems, ensuring that defensive efforts are focused on genuine vulnerabilities rather than measurement artifacts.
Data Availability Statement: All data and materials used in this study are openly available in a public repository: https://github.com/christopherburger/SynEval
Acknowledgments: The authors have no acknowledgments to declare.